Suppose that you generated a forecast with the some data, and you have the following errors:

Week (Et)
1
2 18.2
3 -9.4
4 0.9
5 -4
6 4.6
7

What is the MSE?
Answer #1

Question: Suppose that you generated a forecast with the some data, and you have the following errors:

Answer: MSE = 91.514

Explanation:

Given that:

Error (Et):

Week 2 = 18.2

Week 3 = -9.4

Week 4 = 0.9

Week 5 = -4

Week 6 = 4.6

Now:

To find Mean Squared Error (MSE), first let us square the Error (Et)

That is:

Week 2 = 18.2 = (18.2 x 18.2) = 331.24

Week 3 = -9.4 = (9.4 x 9.4) = 88.36

Week 4 = 0.9 = (0.9 x 0.9) = 0.81

Week 5 = -4 = (4 x 4) = 16

Week 6 = 4.6 = (4.6 x 4.6) = 21.16

Now:

Mean Square Error (MSE) is given by:

MSE = \sum Error^2 / n

Therefore:

MSE = (331.24 + 88.36 + 0.81 + 16 + 21.16) / 5

MSE = 457.57 / 5

MSE = 91.514